The EPA has reported that the average fuel cost for a particular type of car is $800 with a standard deviation of $80. Fuel cost is assumed to be normally distributed. If one of these cars is randomly selected, what is the probability that the fuel cost for this car exceeds $900?

Answer by Boreal(15235)

z=(x-mean)/sd=(900-800)/80=5/4 or 1.25
probability z> 1.25 is 0.1056